How to represent the area of w= x and l = x+2?
How to present the area of (w-3) and l= (x+8)?
Write an open sentence that would indicate that the area of the new and original rectangles are equal.
Write an open sentence that would indicate that the area of the new rectangle is 6 sq. in. more than the area of the original rectangle.

Answer by ewatrrr(24785)

 
Hi,
.
A = Lw = (x+2)x = x^2+2 Original
A = Lw = (x+8)(x-3) = x^2 + 5x-24 New
...
area of the new and original rectangles are equal
x^2+2 = x^2 + 5x-24
...
area of the new rectangle is 6 sq. in. more than the area of the original rectangle.
x^2+2 = x^2 + 5x-18